# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v1

@@ 4.1 High-Level Problem Description
This section provides a high-level overview of the online order fulfillment challenges faced by the industrial partner. Similar to other major e-commerce companies, tens of thousands of orders are placed continuously throughout the day on the website and await processing. Each incoming order is organized by stock-keeping units (SKUs), resulting in one or more order lines. For each order line, there are hundreds of fulfillment centers across the company network eligible to source the SKUs. These centers include primarily physical stores, as well as regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping. The set of available carriers spans from major shipping companies to crowdsourced gig carriers. Figure 1 illustrates the order fulfillment life cycle within the operations of the industrial partner.
@@ 4.2 The Nominal Optimization Problem
This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem, which is a special case of the generic problem described in Section 3.1. A comprehensive list of the notations presented in this section is provided in Appendix for reference.
@@ 4.2.1 Main Decision Variables
Let $\mathcal{I}$ be the set of SKUs and $\mathcal{V}$ be the set of fulfillment options. Here, $\mathcal{V}$ is defined as the Cartesian product $\mathcal{K}\times\mathcal{L}$, where $\mathcal{K}=\{1,\ldots,K\}$ denotes the set of carriers, and =$\mathcal{L}=\{1,\ldots,L\}$ denotes the set of locations.
Each order $o$ needs to solve an independent optimization problem to decide the selections of carrier-location pairs for each SKU. The primary variables are denoted by the vector $\mathbf{z_{o}}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$, where each $z_{ik\ell o}\in\mathbb{Z}$ indicates the quantity of SKU $i$ sourced from location $\ell$ and shipped by carrier $k$ in order $o$.
@@ 4.2.2 Objective Function
The objective of the optimization problem for order $o$ is to minimize the overall fulfillment costs while ensuring timely delivery. The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair. Delivery timeliness is measured by deviations from the desired delivery date, whether early or late. To encourage timeliness, the objective imposes penalties for both early and late deliveries.
The objective function is formally defined as follows. Let $c_{ik\ell o}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$ in order $o$. For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost. For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center. Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
In an ideal setting where all carrier-location deviations are known at the time of fulfillment, let the vector $\mathbf{d}_{o}=(d_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ represent the realized uncertain delivery deviations for order $o$, where each $d_{k\ell o}\in\mathbb{R}$ denotes the number of days that the carrier-location pair $(k,\ell)$ deviates from desired delivery date. A positive deviation indicates late delivery, a zero deviation indicates on-time delivery, and a negative deviation indicates early delivery.
The weighted objective function for order $o$ can be constructed as follows:
$g(\mathbf{z}_{o},\mathbf{d}_{o})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-}]z_{ik\ell o}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell o}\geq 2\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-})z_{ik\ell o}&\text{otherwise},\end{cases}$ (13)
where $d_{k\ell o}^{+}=\max\{0,d_{k\ell o}\}$ and $d_{k\ell o}^{-}=\max\{0,-d_{k\ell o}\}$. This objective function encourages item consolidation within an order by offering a carrier-specific discount when applicable. In addition, it leverages an asymmetric penalty cost function for delivery deviations, with $\gamma^{+}\gg\gamma^{-}$ to primarily discourage late deliveries while also disincentivizing early deliveries to a lesser extent. The objection function is piecewise linear and can be linearized using standard rewritings.
@@ 4.2.3 Operational Constraints
Several operational constraints ensure that the selection of fulfillment center locations and carriers is feasible in practice. For location $\ell\in\mathcal{L}$, let $\operatorname{inv}_{i\ell o}\geq 0$ be the inventory level of SKU $i$ at location $\ell$ and $\operatorname{cap}_{\ell o}\geq 0$ be the available capacity at location $\ell$ when order $o$ is placed. For SKU $i$ in order $o$, let $q_{io}\geq 0$ be the quantity of SKU $i$ in order $o$, and let $e_{ik\ell o}\in\{0,1\}$ be a binary indicator of whether the location-carrier pair $(\ell,k)$ is an eligible fulfillment option.
@@
    $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}z_{ik\ell o}=q_{io},\quad\forall i\in\mathcal{I}$ (14)
@@
    $z_{ik\ell o}\leq q_{io}e_{ik\ell o},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$ (15)
@@
    $\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{inv}_{i\ell o},\quad\forall i\in\mathcal{I},\ell\in\mathcal{L}$ (16)
@@
    $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{cap}_{\ell o},\quad\forall\ell\in\mathcal{L}$ (17)
4.3 Applying the CSO Framework to the Problem
@@
In practice, the delivery deviations of each carrier-location pair cannot be perfectly estimated in advance. To account for this uncertainty for each order $o$, this section models the vector of uncertain delivery deviations $\tilde{\mathbf{d}}_{o}=(\tilde{d}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}\sim\mathbb{P}$ as a $(KL)$-dimensional random variable governed by the probability distribution $\mathbb{P}$. In addition, let $\mathbf{s}_{o}=(\mathbf{s}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ denote the state vector observable for order $o$ before the fulfillment decision is made, where each $\mathbf{s}_{k\ell o}$ is a vector of covariates encompassing information about the fulfillment network status and order details related to carrier-location pair $(k,\ell)$. The conditional distribution of the deviation vector, upon observing $\mathbf{s}_{o}$, is denoted by $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$, and the individual deviation condition distribution is $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ for each $k\in\mathcal{K}$ and $\ell\in\mathcal{L}$.
@@
Applying the CSO framework leads to the following contextual stochastic order fulfillment problem (CSOFP):
(CSOFP) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\Pi_{k\in\mathcal{K},\ell\in\mathcal{L}}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})].$ (18)
Effectively solving Problem (18) requires leveraging data from historical orders. As a specific instance of Problem (2), CSOFP also has an observational dataset. Let $\mathcal{O}$ be the set of historical orders. Let $\mathbf{z}_{o}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$ be the vector of historical decisions made by the industrial partner’s sourcing engine for order $o$. There are numerous missing entries in the observed deviation vector $\mathbf{d}_{o}$ since actual observations depend on the decision vector $\mathbf{z}_{o}$. Let
$\mathbf{D}_{o}:=\{d_{k\ell o}\mid\sum_{i\in\mathcal{I}}z_{ik\ell o}>0,k\in\mathcal{K},\ell\in\mathcal{L}\}$
be the observable costs for order $o$. The dataset available to estimate $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
$\mathbb{D}$ is partitioned into subsets to facilitate learning the individual condition distributions. For carrier $k$ and location $\ell$, the dataset for learning $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ is $\mathbb{D}_{k\ell}=\{(\mathbf{s}_{k\ell m},d_{k\ell m})\}_{m=1}^{M_{k\ell}}$, where $M_{k\ell}$ is the number of times the deviation has been observed for the carrier-location pair $(k,\ell)$ in $\mathbb{D}$.
4.4 Learning the Contextual Distribution of Delivery Deviations
This section describes how to learn Contextual Distribution of Delivery Deviations. It begins by selecting features to construct the covariate vector for each carrier-location pair. To predict the conditional distribution of deviations, two classes of machine learning methods are proposed: probabilistic multi-class classification, which leverages the discrete nature of deviations, and quantile regression, which is known for its effectiveness in distributional prediction (Salari et al. 2022). For simplicity, the index $k\ell$ is omitted in the notations used throughout this section. The data set can then be denoted by $\mathbb{D}=\{(\mathbf{s}_{m},d_{m})\}_{m\in[M]}$.
4.4.1 Contextual Features Selection
Carefully selecting contextual features is crucial for constructing informed covariates for delivery deviation predictions, as various stages in the order life cycle can contribute to deviations. To identify the most important features for predicting deviations, a feature importance analysis was conducted using a tree-based mean regression model. Two methods were examined: Mean Decrease in Impurity (MDI) and Shapley values (SHAP). MDI measures the importance of a feature by calculating the total reduction in impurity (e.g., Gini impurity or entropy) it brings when used in decision trees (Breiman 2001). SHAP values, on the other hand, provide a unified measure of feature importance based on cooperative game theory, attributing to each feature its contribution to the prediction by considering all possible feature combinations (Lundberg et al. 2020).
The features considered are categorized into three main classes, each corresponding to different stages of the order fulfillment process:

### v2

@@ 3.1 High-Level Problem Description
This section provides a high-level overview of the online order fulfillment challenges faced by the industrial partner. Similar to other major e-commerce companies, tens of thousands of orders are placed continuously throughout the day on the website and await processing. Each incoming order is organized by stock-keeping units (SKUs), resulting in one or more order lines. For each order line, there are hundreds of fulfillment centers across the company network eligible to source the SKUs. These centers include primarily physical stores, as well as regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers capable of providing different levels of service, such as Ground, Same-Day, and Overnight shipping. Throughout this paper, “carrier” refers to one of these specific carrier-service pairs. The set of available carriers spans from major shipping companies to crowdsourced gig carriers. Figure 1 illustrates the order fulfillment life cycle within the operations of the industrial partner.
@@ 3.2 The Nominal Optimization Problem
This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem. A comprehensive list of the notations presented in this section is provided in Appendix for reference.
@@ 3.2.1 Main Decision Variables
Let $\mathcal{I}$ be the set of SKUs and $\mathcal{V}$ be the set of fulfillment options. Here, $\mathcal{V}$ is defined as the Cartesian product $\mathcal{K}\times\mathcal{L}$, where $\mathcal{K}$ denotes the set of carriers, and $\mathcal{L}$ denotes the set of locations.
Each order needs to solve an independent optimization problem to decide the selection of carrier-location pairs for each SKU. The primary decision variables are denoted by the vector $\mathbf{z}=(z_{ik\ell})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$, where each $z_{ik\ell}\in\mathbb{Z}$ indicates the quantity of SKU $i$ sourced from location $\ell$ and shipped by carrier $k$.
@@ 3.2.2 Objective Function
The objective of the optimization problem for an order is to minimize the overall fulfillment costs while ensuring timely delivery. The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair. Delivery timeliness is measured by deviations from the desired delivery date, whether early or late. To encourage timeliness, the objective imposes penalties for both early and late deliveries.
The objective function is formally defined as follows. Let $c_{ik\ell}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$. For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost. For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center. Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
In an ideal setting where all carrier-location deviations are known at the time of fulfillment, let the vector $\mathbf{d}=(d_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ represent the realized uncertain delivery deviations, where each $d_{k\ell}$ denotes the number of days that the carrier-location pair $(k,\ell)$ deviates from desired delivery date. A positive deviation indicates late delivery, a zero deviation indicates on-time delivery, and a negative deviation indicates early delivery.
The weighted objective function for an order can be constructed as follows:
$g(\mathbf{z},\mathbf{d})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-}]z_{ik\ell}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell}\geq 2,\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-})z_{ik\ell}&\text{otherwise},\end{cases}$ (1)
where $d_{k\ell}^{+}=\max\{0,d_{k\ell}\}$ and $d_{k\ell}^{-}=\max\{0,-d_{k\ell}\}$. This objective function encourages item consolidation within an order by applying a carrier-specific discount if at least two units are sourced from the same location and shipped by the same carrier. In addition, it leverages an asymmetric penalty cost function for delivery deviations, where $\gamma^{+}\gg\gamma^{-}$. This asymmetry strongly penalizes late deliveries to discourage delays while also applying a smaller penalty to disincentivize early deliveries. Specifically, the penalties can be interpreted as the opportunity cost associated with early deliveries and the lost-sale cost associated with late deliveries. The objective function is piecewise linear and can be linearized using standard rewritings.
@@ 3.2.3 Operational Constraints
Several operational constraints ensure that the selection of fulfillment center locations and carriers is feasible in practice. For location $\ell\in\mathcal{L}$, let $\operatorname{inv}_{i\ell}\geq 0$ be the inventory level of SKU $i$ at location $\ell$ and $\operatorname{cap}_{\ell}\geq 0$ be the available capacity at location $\ell$ when the order is placed. Let $q_{i}\geq 0$ be the quantity of SKU $i$ in the order, and let $e_{ik\ell}\in\{0,1\}$ be a binary indicator of whether the location-carrier pair $(\ell,k)$ is an eligible fulfillment option for SKU $i$. An assignment $\mathbf{z}$ in the feasible set $\mathcal{Z}$ must obey the following constraints.
@@
    $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}z_{ik\ell}=q_{i},\quad\forall i\in\mathcal{I}$ (2)
@@
    $z_{ik\ell}\leq q_{i}e_{ik\ell},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$ (3)
@@
    $\sum_{k\in\mathcal{K}}z_{ik\ell}\leq\operatorname{inv}_{i\ell},\quad\forall i\in\mathcal{I},\ell\in\mathcal{L}$ (4)
@@
    $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}z_{ik\ell}\leq\operatorname{cap}_{\ell},\quad\forall\ell\in\mathcal{L}$ (5)
3.3 The Contextual Stochastic Order Fulfillment Problem
In practice, the delivery deviations of each carrier-location pair cannot be perfectly estimated in advance. To account for this uncertainty for each order, this section models the vector of uncertain delivery deviations $\tilde{\mathbf{d}}=(\tilde{d}_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}\sim\mathbb{P}$ as a $(|\mathcal{K}||\mathcal{L}|)$-dimensional random vector governed by a joint probability distribution $\mathbb{P}$. Since deviations measured in days only take on discrete values in the dataset, each $\tilde{d}_{k\ell}$ inherently follows a discrete probability distribution. It is further assumed that every $\tilde{d}_{k\ell}$ shares the same finite, ordered support. To formalize this, let $C$ be the total number of distinct deviation values and denote them by $\xi_{1}<\xi_{2}<\ldots<\xi_{C}$. By definition, any realization $d_{k\ell}$ of $\tilde{d}_{k\ell}$ must lie in this set: $d_{k\ell}\in\{\xi_{1},\xi_{2},\ldots,\xi_{C}\}.$
@@ 3.3 The Contextual Stochastic Order Fulfillment Problem
In addition, let $\mathbf{s}=(\mathbf{s}_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ denote the state vector observable to the order before the fulfillment decision is made, where each $\mathbf{s}_{k\ell}$ is a vector of covariates encompassing information about order details and fulfillment network status related to carrier-location pair $(k,\ell)$. The contextual distribution of the deviation vector, upon observing $\mathbf{s}$, is denoted by $\mathbb{P}(\mathbf{d}\mid\mathbf{s})$, and the individual deviation contextual distribution is $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$ for each $k\in\mathcal{K}$ and $\ell\in\mathcal{L}$.
@@
This leads to the following contextual stochastic order fulfillment problem (CSOFP):
(CSOFP) $\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (6)
Solving Problem (6) in practice requires leveraging data from past orders. Unlike classical CSO formulations where the entire random cost vector is observed, CSOFP is based on an observational dataset of past decisions and their realized costs. Let $\mathcal{O}$ be the set of historical orders. For each $o\in\mathcal{O},$ define the decision vector $\mathbf{z}_{o}=(z_{ik\ell,o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}.$ Since deviations were observed only for chosen $(i,k,\ell)$ triples, the full deviation vector $\mathbf{d}_{o}=(d_{k\ell,o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ contains many unobserved entries. The subset of realized deviations for order $o$ is:
$\mathbf{D}_{o}:=\{d_{k\ell,o}\mid\sum_{i\in\mathcal{I}}z_{ik\ell,o}>0,k\in\mathcal{K},\ell\in\mathcal{L}\}.$
The resulting dataset available for estimating the contextual distribution $\mathbb{P}(\mathbf{d}\mid\mathbf{s})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
4 Methodology: A Generic CSO Framework for Observational Data
Solving CSOFP under observational data requires: (1) learning a contextual distributional oracle from partial realizations, and (2) optimizing with that oracle in a tractable way.
This section lays out a generic contextual stochastic optimization (CSO) framework that does exactly that. In Section 4.1, the abstract CSO problem (denoted by ${\cal P}$) is defined. Section 4.2 shows how to construct a contextual distribution oracle $\mathcal{M}(\mathbf{s})$ from partial observations. Sections 4.3 and 4.4 introduce two solution paradigms—Contextual Sample Average Approximation and Contextual Robust Optimization—which use $\mathcal{M}(\mathbf{s})$ as input. Finally, Section 4.6 maps all of the above back to the omnichannel multi-courier CSOFP introduced in Section 3.
4.1 Generic Contextual Stochastic Optimization (CSO) Problem
Problem ${\cal P}$ consists of fulfilling a set of incoming requests sequentially, each of which specifies demands for a subset of products. Let $\mathcal{I}$ be the set of products, $\mathcal{V}$ the set of selection options (e.g., sourcing or transport modes), and $\mathcal{Z}\subseteq\mathbb{Z}_{+}^{|\mathcal{I}|\times|\mathcal{V}|}$ the feasible assignment set. Each request arrives with known demand quantities $\{q_{i}:i\in\mathcal{I}\}$ at decision time. The decision variable $\mathbf{z}=(z_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\in\mathcal{Z}$ assigns quantity $z_{iv}$ of product $i$ to option $v$.
Let $\tilde{\mathbf{c}}=(\tilde{c}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\sim\mathbb{P}$ denote the random cost vector. Assume that $\tilde{\mathbf{c}}$ is element-wise positive and the cost function $g(\mathbf{z},\tilde{\mathbf{c}})$ is piecewise linear in $\mathbf{z}$. Introduce, for $j=0,1,\dots,n$, the breakpoint vectors
[…]
so that each segment is defined as
$[\mathbf{b}^{j-1},\,\mathbf{b}^{j}]=\bigl{\{}\mathbf{z}:b^{\,j-1}_{iv}\leq z_{iv}\leq b^{j}_{iv}\;\forall\,i\in\mathcal{I},\;v\in\mathcal{V}\bigr{\}}.$
Let $\mathbf{a}^{j}=(a^{j}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}$ be the marginal‐cost adjustment on segment $j$. Then
$g(\mathbf{z},\tilde{\mathbf{c}})=(\tilde{\mathbf{c}}+\mathbf{a}^{j})^{\mathsf{T}}\mathbf{z}\quad\text{if}\quad\mathbf{z}\in[\mathbf{b}^{j-1},\mathbf{b}^{j}],\;j=1,\dots,n,$ (7)
thereby capturing volume‐dependent pricing behaviors.
For each request, the fulfillment decision solves the risk‐neutral stochastic optimization problem
$\min_{\mathbf{z}\in\mathcal{Z}}\;\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}}\bigl{[}g(\mathbf{z},\tilde{\mathbf{c}})\bigr{]}.$ (8)
The piecewise‐linear cost in (7) partitions the decision space into shipment‐size tiers determined by the breakpoints $\{\mathbf{b}^{j}\}$. Within tier $j$, the marginal cost is $\tilde{\mathbf{c}}+\mathbf{a}^{j}$, which shifts to $\tilde{\mathbf{c}}+\mathbf{a}^{j+1}$ upon entering tier $j+1$. This formulation accommodates economies of scale (declining unit costs at higher volumes), quantity discounts (step‐down pricing beyond thresholds), capacity surcharges (unit‐cost premiums when exceeding lower‐cost bands), and other tiered‐pricing schemes.
In Problem (8), the true distribution $\mathbb{P}$ is not readily available. However, side information often exists for each product-option pair. Let $\mathbf{s}_{iv}$ denote the covariate vector for product $i$ under selection option $v$ (e.g., network state, weather, current demand). All these covariate vectors are aggregated into the state vector $\mathbf{s}=(\mathbf{s}_{iv})_{i\in\mathcal{I},v\in\mathcal{V}}$, which encompasses all contextual information available at the time of decision. The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$, yielding the Contextual Stochastic Optimization (CSO) problem
$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (9)
Let $\tilde{\mathbf{c}}_{v}=(\tilde{c}_{iv})_{i\in\mathcal{I}}$ and $\mathbf{s}_{v}=(\mathbf{s}_{iv})_{i\in\mathcal{I}}$ for each $v\in\mathcal{V}$. Under the assumption that, conditional on its own covariates, each component $c_{v}$ is independent across options $v\in\mathcal{V}$, one has
$\mathbb{P}(\mathbf{c}\mid\mathbf{s})=\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v}),$
where $\mathbb{P}_{v}(\cdot\mid\mathbf{s}_{v})$ is the contextual distribution for option $v$. This assumption is reasonable when each option’s cost depends primarily on its own covariates. Shared features, such as weather, create similarity but do not induce direct dependence. Consequently, Problem (9) can be written as
$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (10)
4.2 Constructing a Contextual Distribution Oracle from Observational Data
The contextual distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ is also generally unknown but can often be inferred from historical observations. In many real-world COS applications (including CSOFP) however, data are observational, i.e., only the costs corresponding to decisions actually taken were recorded. As a result, the dataset available to estimate $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ takes the form $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{C}_{o})\}_{o=1}^{|\mathbb{D}|}$, where $\mathbf{s}_{o}=(s_{iv,o})_{i\in\mathcal{I},\;v\in\mathcal{V}}~{}\text{and}~{}\mathbf{C}_{o}=(\mathbf{c}_{v,o})_{v\in\mathcal{V}}$ collect the context and realized costs, with $\mathbf{c}_{v,o}$ observed only if option $v$ was selected in request $o$.
The learning task consists in finding a full‑vector forecasting oracle $\mathcal{M}$ that approximates $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ in a supervised manner using the dataset $\mathbb{D}$: $\mathcal{M}$ receives a context $\mathbf{s}$ as input and returns a probability distribution $\mathcal{M}(\mathbf{s})$ on the cost vector $\tilde{\mathbf{c}}$. Under the conditional independence assumption from the previous section, learning $\mathbb{P}(\mathbf{c}\mid\mathbf{s})$ amounts to learning $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})$ for each selection option $v$ separately. Partition the full dataset $\mathbb{D}$ into subsets $\mathbb{D}_{v}=\{(\mathbf{s}_{v,1},\mathbf{c}_{v,1}),(\mathbf{s}_{v,2},\mathbf{c}_{v,2}),\ldots\},$ representing the dataset for option $v$. A machine learning model $\mathcal{M}_{v}$ is then trained on $\mathbb{D}_{v}$ to approximate each $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})$. Assemble the oracle as the product $\mathcal{M}(\mathbf{s})=\bigl{(}\mathcal{M}_{v}(\mathbf{s}_{v})\bigr{)}_{v\in\mathcal{V}},$ which generates joint forecasts by treating each component independently. Note that if every $\mathbf{s}_{v}$ shares the same structure and dimension, a single learning architecture can serve as $\mathcal{M}_{v}$ for all $v\in\mathcal{V}$, trained on the union $\cup_{v\in\mathcal{V}}\mathbb{D}_{v}$.
By reducing the high‑dimensional joint learning task to $|\mathcal{V}|$ manageable subproblems, this decomposition turns observational data into a black‑box contextual distribution oracle $\mathcal{M}(\mathbf{s})$, which will serve as the input for the solution methods developed in Section 4.3 (C-SAA) and Section 4.4 (C-RO).
4.3 Contextual Sample Average Approximation (C-SAA)
Problem (10) is approximated by generalizing the Sample Average Approximation (SAA) method (Kleywegt et al. 2002) to contextual stochastic optimization. Given a request with covariates $\mathbf{s}$, the Contextual Sample Average Approximation (C-SAA) proceeds in two stages: